A weight if 20N hangs from a fixed point by a light inextensible string. It is pulled aside by a horizontal force with the string inclined at an angle of 30° to the vertical. What is the tension in the string?
1
Expert's answer
2019-04-04T09:03:38-0400



The weight will be in balance when:

"\begin {cases}
T \sin \alpha = F \\ 
T \cos \alpha = mg
\end{cases}" Then "T = \frac{mg}{cos \alpha} = \frac{20N}{\cos (30^\circ)}" , so


T = 23.09N